If an ideal gas law has apressure of 7.19atm a temperature of 99.8 celcius a volume of 64.87centimetre cube. find the number of moles


1
Expert's answer
2022-02-14T21:33:03-0500

PV= nRT

"PV= nRT"

7.19× 64.87× 10^-6= n× 8.314 × 372.95

n = 0.047/3100.7

n = 1.507× 10^-7 moles